# Learning Mixtures of Linear Dynamical Systems (MoLDS) via Hybrid Tensor–EM Method

### v1

@@ 3.3 Tensor Initialization for MoLDS
The tensor initialization leverages the key insight that MoLDS can be reformulated as MLR through lagged input representations. This transformation exposes the mixture structure in high-order moments, enabling algebraic recovery of component parameters via tensor decomposition (see Appendix A for details). The method works by exploiting the impulse-response (Markov parameter) representation of LDSs. We construct second- and third-order cross moments of the lagged input–output data, denoted $M_{2}$ and $M_{3}$ (see Appendix A.2). Then we apply whitening transformations and decompose the resulting tensor into weights and Markov parameter estimates. At last, we recover LDS state-space parameters through the Ho-Kalman realization Oymak & Ozay (2019). The core mathematical insight is that if $M_{2}$ and $M_{3}$ are appropriately constructed from sub-sampled lagged covariates, the whitened tensor
@@
is symmetric and orthogonally decomposable, where $\alpha_{k}\in\mathbb{R}^{K}$ are the whitened representations of the regression vectors (scaled Markov parameters) for each LDS component, and $p_{k}$ are the corresponding mixture weights. This decomposition is unique and recovers the mixture components $\{\alpha_{k},p_{k}\}$ up to permutation and sign. Importantly, we employ SMD for this decomposition, which is empirically shown to be more stable in Section 4.1. We then apply the Ho-Kalman realization Oymak & Ozay (2019) to recover the state-space parameters for each component, which are then used as principled initializations for the subsequent refinement stage. The complete procedure is provided in Algorithm 3 in Appendix A.4, together with the MoLDS-to-MLR reformulation and tensor construction details in Appendix A.
Figure 1: Comparison of SMD-Tensor and RTPM-Tensor methods for MoLDS. (a,b) Mean Markov parameter estimation errors for the SMD method decrease as the number of trajectories as NN and/or $T$ increase; (c) (c) Heatmap of the best trial result across all $(N,T)$ configurations. (d–f) Corresponding results for the RTPM method.(g) Difference in mean Markov errors between RTPM and SMD (positive values indicate SMD performs better). (h) Scatter plot comparing mean Markov errors of RTPM vs. SMD across configurations, with SMD outperforming in $91\%$ of cases. (i) Example recovery for $N=T=1280$, where SMD recovers both mixture weights and Markov parameters more accurately.
@@ 3.4 EM Refinement for MoLDS
The tensor initialization provides globally consistent estimates of mixture weights and system matrices, but does not recover the noise parameters $(Q_{k},R_{k})$ nor achieve optimal statistical accuracy. We therefore need to refine these estimates using a full Kalman EM algorithm that maximizes the observed-data likelihood.
Our EM formulation extends classical EM to the MoLDS setting by computing trajectory-wise responsibilities via Kalman filter likelihoods, then updating parameters from responsibility-weighted sufficient statistics (see Appendix B and Algorithm 4 for details). In brief, at iteration $t$, given current parameters $\hat{\theta}^{(t)}=\{(\hat{p}_{k},\hat{A}_{k},\hat{B}_{k},\hat{C}_{k},\hat{D}_{k},\hat{Q}_{k},\hat{R}_{k})\}_{k=1}^{K}$, the E-step computes responsibilities
$\gamma_{i,k}^{(t)}\;=\;\frac{\exp\!\big(\log\hat{p}_{k}^{(t)}+\log p(y_{i,0:T_{i}-1}\mid u_{i,0:T_{i}-1},\hat{\theta}_{k}^{(t)})\big)}{\sum_{r=1}^{K}\exp\!\big(\log\hat{p}_{r}^{(t)}+\log p(y_{i,0:T_{i}-1}\mid u_{i,0:T_{i}-1},\hat{\theta}_{r}^{(t)})\big)}.$ (4)
Next, we use a Kalman smoother to compute responsibility-weighted sufficient statistics $S_{k}^{(t)}$ for each component. The M-step then updates all parameters via closed-form maximum likelihood estimates
@@ 4.1 SMD-tensor method provides more reliable recovery on simulated MoLDS
We demonstrate that the employed SMD–Tensor method reliably recovers parameters in synthetic MoLDS with a small number of components, low latent dimensionality, and well-separated dynamics. We empirically compare SMD–Tensor against the RTPM-Tensor baseline across a wide range of such settings. Our results show that SMD-Tensor consistently outperforms RTPM–Tensor on multiple metrics. In particular, when the number of mixture components is small and sufficient data are available, SMD achieves near-optimal recovery of both the mixture weights and the Markov parameters.
@@
The first row of Figure 1 reports results of the SMD-Tensor method for a $K=3$ mixture model with LDS dimensions $n=2,m=p=1$. The LDS parameters are randomly generated, with eigenvalues of $A$ constrained to lie inside the unit circle. We vary the trajectory length $T$ and the number of trajectories $N$, and for each $(N,T)$ configuration. We run multiple independent trials, calculating the discrepancy between estimated and true Markov parameters. The second row of Figure 1 shows the result for RTPM-Tensor. It is noticed that the mean Markov parameter errors decrease with increasing trajectory length and number of trajectories for both methods, but the trend is more pronounced and stable for SMD, as reflected in the heat maps (Figure 1(c,f)). Across nearly all $(N,T)$ cases, SMD yields consistently lower errors (Figures 1(g,h)). In cases with larger $T$ and $N$, SMD achieves highly accurate recovery of both mixture weights and Markov parameters (Figure 1(c1)), underscoring its robustness and reliability for more complex conditions.
4.2 Tensor–EM improves robustness and accuracy for complex synthetic MoLDS
@@
In complex MoLDS settings with many components, purely tensor-based methods and randomly initialized EM often fail to achieve accurate recovery, either due to noisy parameter estimates or convergence to poor local optima. Our proposed Tensor–EM approach overcomes these limitations by combining globally consistent tensor initialization with EM-based refinement, resulting in more robust and accurate learning. We demonstrate these advantages in this section.
@@
Figure 2 presents results on a simulated $K=6$ MoLDS ($n=3$, $m=p=2$) with zero direct feedthrough ($D=0$). Across metrics including Markov parameter errors, mixture weight errors, aggregate LDS parameter errors, and iteration counts, Tensor-EM consistently outperforms all baselines. It achieves substantially lower Markov and weight errors (panels b and c), while maintaining an aggregate parameter error comparable to the pure tensor solution (panel d). [^2] The aggregate parameter error, i.e., the geometric mean of $A,B,C$ errors, is a coarse summary: unlike Markov parameter errors, it may be less precise since $A,B,C$ can differ by similarity transformations without altering the underlying dynamics. At the same time, Tensor-EM converges in much fewer iterations than random EM (e), leading to a better error-efficiency tradeoff (f). The radar plot (a) summarizes these improvements, showing that Tensor-EM achieves the most balanced performance across all evaluation metrics. These results highlight that Tensor-EM effectively combines the strengths of tensor methods and EM and enable more reliable parameter recovery in synthetic MoLDS settings where standalone tensor or randomly initialized EM approaches often underperform.
@@
Figure 2: Performance comparison of pure tensor, Tensor-EM, and random-initialized EM on a simulated MoLDS. (a) Relative accuracy radar plot across metrics of Markov parameter accuracy, weight accuracy, and aggregate parameter accuracy. (b-c) Tensor-EM achieves the lowest Markov parameter and weight errors, while random EM performs the worst and shows high variability. (d) reports aggregate parameter errors (geometric mean of $A,B,C$ errors). (e) Tensor–EM converges in far fewer iterations than random EM, highlighting efficiency. (f) Error-efficiency plot shows that Tensor-EM combines low error with moderate iteration cost, yielding robust and accurate recovery.
@@
5 Tensor–EM MoLDS Provides Reliable and Interpretable Recovery on Real-world Applications
@@
We next apply the proposed MoLDS method to two real-world neural datasets. Figure 3 gives an overview of the approach and examples from these two datasets.
Figure 3: Overview of MoLDS and its applications: (a) Schematic of the MoLDS model structure; MoLDS models observed trajectories as emanating from a set of LDS’s. (b) The Area2 Dataset contains neural trajectories from monkey primary somatosensory cortex; reach trajectories and firing rate from an example neuron. (c) The PMd Dataset contains recordings from monkey dorsal premotor cortex; hand speed and neurons’ rasters.
@@ 5.1 Area2 dataset
We first employ a neural dataset from a single recording session of a macaque performing center-out reaches, with neural activity recorded from the somatosensory cortex. During the experiment, the monkey used a manipulandum to direct a cursor from the center of a display to one of eight targets, and neural activity was recorded from Brodmann’s area2 of the somatosensory cortex (see Chowdhury et al. (2020) for experimental details, and Miller (2022) for dataset location). The neural recordings consist of $65$ single-unit spike times, which are converted to spike rates. In addition to neural data, the position of the monkey’s hand, cursor position, force applied to the manipulandum, and hand velocity were recorded during the experiment. For our MoLDS setting, we take the hand velocity variables as inputs, and the $65$-dim observations are reduced to $6$-dim and are regarded as outputs for simplicity by using the standard PCA. There are $8$ directions in the task (as shown in Figure 3 (b)), and each direction has multiple trials of input-output trajectories.
@@
We evaluate the MoLDS with Tensor–EM and Random-EM methods using a standard train/validation/test split of the dataset (see full pipeline in App. D). For each hyperparameter setting (including $K$), models are trained on the training set and scored on the validation set using negative log-likelihood (NLL), one-step-ahead RMSE, and BIC. The model used for test-time analysis is the one minimizing validation BIC (we also report NLL/RMSE). For trial $i$ and component $k$, we compute a responsibility $r_{ik}\propto\exp(\ell_{k}^{(i)})$, where $\ell_{k}^{(i)}$ is the one-step Kalman log-likelihood under component $k$. For each movement direction, the dominant component is the $\arg\max_{k}$ of the mean responsibility across its trials.
@@
Figure 4 summarizes results. As shown in Fig. 4(a), validation criteria consistently favor a $3$-component MoLDS trained with Tensor-EM. (b) Test one-step predictions $\hat{y}_{t}$ closely track observations $y$. (c) Dominant-component maps reveal three cross-trial clusters aligned with direction; (e) usage fractions quantify prevalence on the test set.
@@
Figure 4: MoLDS application on Area2 Dataset: (a) Validation metrics for different $K$. (b) One-step predictions for two example trials using corresponding LDS components of the MoLDS selected by the lowest validation metrics. (c) Agreement between MoLDS trial assignments (outer ring) and per-direction single-LDS clusters (inner ring); the SLDS-based method (middle ring) cannot provide correct cross-trial clusters. (d) Per-direction impulse response (Markov parameter) curves for assigned MoLDS component vs. the corresponding single-LDS (titles report cosine similarity). (e) MoLDS component usage fractions on held-out test trials.
@@
In this setup, we also compare Tensor-EM method with the supervised learning results of LDS, where we train one LDS on all trials under each direction. The per-direction LDS baseline closely matches MoLDS-both in impulse responses (high cosine similarity) and in trial groupings (Fig. 4c-d). We also train an SLDS in a similar unsupervised way of MoLDS (no direction labels) and find that it does not yield meaningful cross-trial clusters here, while it is effective for within-trial regime switches (see Figure 6 in Appendix). These highlight MoLDS’s strength in capturing between-trial heterogeneity. Importantly, when compared with the Random-EM method, Tensor-EM offers a key advantage: the tensor initialization step yields a stable starting point that reduces variability across runs, leading to more consistent parameter recovery and model structure as seen in Figure 4(a).
@@ 5.2 PMd Dataset
We next evaluate our method on recordings from monkey dorsal premotor cortex (PMd) during more complex reaches, in which trial-wise movement directions are continuously distributed over the circle. Full experimental details are in Lawlor et al. (2018b) and the dataset is provided in Perich et al. (2018). The trials’ reach direction spans the full polar range rather than discrete directions as in the Area2 dataset. We analyze PMd activity with the $5$-dimensional kinematic as inputs, i.e., $(\text{x\mbox{-}velocity},\text{y\mbox{-}velocity},\text{x\mbox{-}acceleration},\text{y\mbox{-}acceleration},\text{speed})$.
Figure 5: MoLDS application on PMd Dataset: (a) Validation metrics for different $K$ (NLL, RMSE, BIC). (b) Test one-step predictions from the validation-chosen MoLDS. (c) Angle bins: dominant component; PD arrows (responsibility-weighted). (d) Impulse response magnitude of MoLDS components. (e) MoLDS component usage fractions on held-out test trials.
Following the Area2 protocol, we train MoLDS across hyperparameters on the PMd training split and select the model by validation criteria (NLL/RMSE/BIC). Figure 5(a) reports validation results, and Fig. 5(b) shows test one-step-ahead reconstructions.
To accommodate continuously distributed reach angles, trials are uniformly binned on the circle (12 bins). Within each bin, the dominant component is the one with the largest mean responsibility. Figure 5(c) colors bins by their dominant LDS component and overlays responsibility-weighted preferred-direction (PD) arrows for each component. Figure 5(d) plots impulse-response magnitudes versus lag of each component, revealing component-specific gains and temporal decay (Appendix D.2, Fig. 7, further decomposes input-channel responses). Component usage fractions on held-out test trials are shown in Fig. 5(e). Overall, the mixture components specialize in distinct movement directions and exhibit distinct dynamical response profiles.

### v2

@@ 3.3 Tensor Initialization for MoLDS
The tensor initialization leverages the key insight that MoLDS can be reformulated as MLR through lagged input representations (Rui & Dahleh, 2025). This transformation exposes the mixture structure in high-order moments, enabling algebraic recovery of component parameters via tensor decomposition (see Appendix B for details). The method works by exploiting the impulse-response (Markov parameter) representation of LDSs. We first construct second- and third-order cross moments of the lagged input-output data, denoted as $M_{2}$ and $M_{3}$
where $\mathcal{N}_{2}$ and $\mathcal{N}_{3}$ are two disjoint subsets of samples, $\tilde{y}_{j}=y_{i,t}$ is the observed output, $v_{j}$ is the normalized lagged input, $I_{d}$ and $\mathcal{E}$ are correction terms (see Appendix B.2 for notations). Then we apply the whitening transformation $W$ and decompose the resulting tensor into weights and Markov parameter estimates. At last, we recover LDS state-space parameters through the Ho-Kalman realization (Oymak & Ozay, 2019). The core mathematical insight is that if $M_{2}$ and $M_{3}$ are appropriately constructed from sub-sampled lagged covariates, the whitened tensor
@@
is symmetric and orthogonally decomposable, where $\alpha_{k}\in\mathbb{R}^{K}$ are the whitened representations of the regression vectors (scaled Markov parameters) for each LDS component, and $p_{k}$ are the corresponding mixture weights. This decomposition is unique and recovers the mixture components $\{\alpha_{k},p_{k}\}$ up to permutation and sign. Importantly, we employ SMD (Kuleshov et al., 2015) for this decomposition step, which is shown to be more stable and accurate empirically in Section 4.1. We then apply the Ho-Kalman realization (Oymak & Ozay, 2019) to recover the state-space parameters for each component, which are then used as principled initializations for the subsequent refinement stage. The complete procedure is provided in Algorithm 3 in Appendix B.4, together with the MoLDS-to-MLR reformulation and tensor construction details in Appendix B.
@@ 3.4 EM Refinement for MoLDS
The tensor initialization provides globally consistent estimates of mixture weights and system matrices, but does not recover the noise parameters $(Q_{k},R_{k})$ nor achieve optimal statistical accuracy. We therefore need to refine these estimates using a full Kalman filter-smoother EM algorithm that maximizes the observed-data likelihood.
Our EM formulation extends classical mixture EM to the MoLDS setting by computing trajectory-wise responsibilities via Kalman filter likelihoods, then updating parameters from responsibility-weighted sufficient statistics (see Appendix C and Algorithm 4 for details). In brief, at iteration $t$, given current parameters $\hat{\theta}^{(t)}=\{(\hat{p}_{k}^{(t)},\hat{A}_{k}^{(t)},\hat{B}_{k}^{(t)},\hat{C}_{k}^{(t)},\hat{D}_{k}^{(t)},\hat{Q}_{k}^{(t)},\hat{R}_{k}^{(t)})\}_{k=1}^{K}$, the E-step computes responsibilities
$\gamma_{i,k}^{(t)}\;=\;\frac{\exp\!\big(\log\hat{p}_{k}^{(t)}+\log p(y_{i,0:T_{i}-1}\mid u_{i,0:T_{i}-1},\hat{\theta}_{k}^{(t)})\big)}{\sum_{k=1}^{K}\exp\!\big(\log\hat{p}_{k}^{(t)}+\log p(y_{i,0:T_{i}-1}\mid u_{i,0:T_{i}-1},\hat{\theta}_{k}^{(t)})\big)}.$ (5)
Next, we use a Kalman smoother to compute responsibility-weighted sufficient statistics $S_{k}^{(t)}$ for each component. The M-step then updates all parameters via closed-form maximum likelihood estimates (MLE)
@@ 4.1 SMD-tensor method provides more reliable recovery on simulated MoLDS
We demonstrate that the employed SMD-Tensor method reliably recovers parameters in synthetic MoLDS with a small number of components, low latent dimensionality, and well-separated dynamics. We empirically compare SMD-Tensor against the RTPM-Tensor baseline across a wide range of such settings. Our results show that SMD-Tensor consistently outperforms RTPM-Tensor on multiple metrics. In particular, when the number of mixture components is small and sufficient data are available, SMD achieves near-optimal recovery of both the mixture weights and the Markov parameters.
Figure 2: Comparison of SMD-Tensor and RTPM-Tensor methods for MoLDS: (a,b) Mean Markov parameter estimation errors of the SMD method decrease as the number of trajectories as NN and/or $T$ increase; (c) Heatmap of the best trial result across all $(N,T)$ configurations. (d-f) Corresponding results for the RTPM method. (g) Scatter plot comparing mean Markov errors of RTPM vs. SMD across configurations, with SMD outperforming in $91\%$ of cases. (h) Difference in mean Markov errors between RTPM and SMD (positive values indicate SMD performs better). (i) Example recovery for $N=T=1280$, where SMD recovers both mixture weights and Markov parameters more accurately.
@@
The first row of Figure 2 reports results of the SMD-Tensor method for a $K=3$ mixture model with LDS dimensions $n=2,m=p=1$. The LDS parameters are randomly generated, with eigenvalues of $A$ constrained to lie inside the unit circle. We vary the trajectory length $T$ and the number of trajectories $N$ for each configuration. We run multiple independent trials, calculating the discrepancy between estimated and true Markov parameters. The second row of Figure 2 shows the result for the RTPM-Tensor method. It is noticed that the mean Markov parameter errors decrease with increasing trajectory length and number of trajectories for both methods, but the trend is more pronounced and stable for SMD, as reflected in the heat maps (Figure 2c,f). Across nearly all $(N,T)$ cases, SMD yields consistently lower errors (Figures 2g,h). In cases with larger $T$ and $N$, SMD achieves highly accurate recovery of both mixture weights and Markov parameters (Figure 2i), underscoring its robustness and reliability for the MoLDS setting.
4.2 Tensor-EM improves robustness and accuracy for complex synthetic MoLDS
@@
In complex MoLDS settings with many components, purely tensor-based methods and randomly initialized EM often fail to achieve accurate recovery, either due to noisy parameter estimates or convergence to poor local optima. Our proposed Tensor-EM approach overcomes these limitations by combining globally consistent tensor initialization with EM-based refinement, resulting in more robust and accurate learning. We demonstrate these advantages in this section.
@@
Figure 3 presents results on a simulated $K=6$ MoLDS ($n=3$, $m=p=2$) with zero direct feedthrough ($D=0$). Across metrics including Markov parameter errors, mixture weight errors, aggregate LDS parameter errors, and iteration counts, Tensor-EM consistently outperforms all baselines. It achieves substantially lower Markov and weight errors (panels b and c), while maintaining an aggregate parameter error comparable to the pure tensor solution (panel d). [^2] The aggregate parameter error, i.e., the geometric mean of $A,B,C$ errors, is a coarse summary: unlike Markov parameter errors, it may be less precise since $A,B,C$ can differ by similarity transformations without altering the underlying dynamics. At the same time, Tensor-EM converges in much fewer iterations than random EM (panel e), leading to a better error-efficiency tradeoff (panel f). The radar plot ( Figure 3a) summarizes these improvements, showing that Tensor-EM achieves the most balanced performance across all evaluation metrics. These results highlight that Tensor-EM effectively combines the strengths of tensor methods and EM and enables more reliable parameter recovery in synthetic MoLDS settings where standalone tensor and randomly initialized EM approaches often underperform.
@@
Figure 3: Performance comparison of pure tensor, Tensor-EM, and random-initialized EM on simulated MoLDS: (a) Relative accuracy radar plot across metrics of Markov parameter accuracy, weight accuracy, and aggregate parameter accuracy. (b-c) Tensor-EM achieves the lowest Markov parameter and weight errors, while random EM performs the worst and shows high variability. (d) reports aggregate parameter errors (geometric mean of $A,B,C$ errors). (e) Tensor-EM converges in far fewer iterations than random EM, highlighting its efficiency. (f) The error-efficiency plot shows that Tensor-EM combines low error with moderate iteration cost, yielding robust and accurate recovery.
@@
5 Tensor-EM MoLDS Provides Reliable and Interpretable Recovery on Real-world Applications
@@
We next apply the proposed MoLDS method to two real-world neural datasets.
@@ 5.1 Area2 dataset
Figure 4: Neural datasets overview: (a) The Area2 Dataset contains neural trajectories from monkey primary somatosensory cortex; reach trajectories and firing rate from multiple neurons. (b) The PMd Dataset contains recordings from monkey dorsal premotor cortex; hand speed and neurons’ rasters.
@@
We first analyze a neural dataset from a single recording session of a macaque performing center-out reaches, with neural activity recorded from the somatosensory cortex (Figure 4a). During the experiment, the monkey used a manipulandum to direct a cursor from the center of a display to one of eight targets, and neural activity was recorded from Brodmann’s area2 of the somatosensory cortex (see (Chowdhury et al., 2020) for experimental details, and (Miller, 2022) for dataset location). The neural recordings consist of $65$ single-unit spike times, which are converted to spike rates. In addition to neural data, the position of the monkey’s hand, cursor position, force applied to the manipulandum, and hand velocity were recorded during the experiment. There are $8$ directions in the task, and each direction has multiple trials of trajectories. For the MoLDS fitting, we extract the movement-related segment of each trial, defined as the window from $100$ ms before to $500$ ms after movement onset. The $65$-dimensional observations are reduced to $6$-dimension by using the standard PCA, which explains more than $90$ percent variance of neural activities. These PCs are then taken as outputs, while the hand velocity variables are taken as inputs. In addition, we also evaluate the dataset with $20$ PCs using MoLDS, and consistent results are found (see Figure 11), which confirms the effectiveness of the Tensor-EM algorithm to fit a MoLDS model on this dataset.
@@
We evaluate the MoLDS with Tensor-EM, Random-EM, and pure tensor methods using a standard train/validation/test split of the dataset (see full pipeline in App. E). For each hyperparameter setting (including $K$), models are trained on the training set and scored on the validation set using negative log-likelihood (NLL), one-step-ahead RMSE, and BIC. The model used for test-time analysis is the one minimizing validation BIC (we also report NLL/RMSE). For trial $i$ and component $k$, we compute a responsibility $r_{ik}\propto\exp(\ell_{k}^{(i)})$, where $\ell_{k}^{(i)}$ is the one-step Kalman log-likelihood under component $k$. For each movement direction, the dominant component is the $\arg\max_{k}$ of the mean responsibility across its trials.
@@
Figure 5 summarizes the results. As shown in Figure 5a, the validation criteria consistently favor a $3$-component MoLDS trained with Tensor-EM. Moreover, the one-step predictions $\hat{y}_{t}$ closely track observations $y$ on the test data as in Figure 5b. In Figure 5c, the dominant-component maps reveal three cross-trial clusters aligned with directions, and the usage fractions in Figure 5d quantify the prevalence of MoLDS components on the test set.
@@
Figure 5: MoLDS application on Area2 Dataset: (a) Validation metrics of MoLDS for different $K$. (b) One-step predictions for two example trials using corresponding LDS components from the MoLDS selected by the lowest validation metrics. (c) Agreement between Tensor-EM MoLDS trial assignments (outer ring) and per-direction LDS, SLDS, single-trial LDS, Random-EM MoLDS clusters (inner rings in each polar plot); the SLDS-based method cannot provide meaningful cross-trial clusters. (d) Tensor-EM MoLDS component usage fractions on held-out test trials.
@@
In this setup, we also compare the Tensor-EM MoLDS method with the supervised learning results of LDS, where we train (1) one LDS on all trials where the monkey reaches in a specific direction (per-dir LDS), and separately (2) a separate LDS fit on each trial regardless of reaching direction (single-trial LDS). Finally, we cluster the parameters of the different LDS’s (per-dir or single-trial). See Figure 8 in the Appendix for a representation of the clustered parameters and more details. The per-direction LDS baseline closely matches the result of MoLDS with the Tensor-EM method rather than Random-EM in trial groupings (see Figure 5c), and the impulse responses are highly similar (see Figure 10 in the Appendix). We also train an SLDS in similar unsupervised way as MoLDS (no direction labels) and find that it does not yield meaningful cross-trial clusters here, while it is effective for possible within-trial regime switches (see Figure 9 in Appendix). These highlight MoLDS’s strength in capturing between-trial heterogeneity. In addition, when compared with the Random-EM method, Tensor-EM offers a key advantage: the tensor initialization step yields a stable starting point that reduces variability across runs, leading to more consistent parameter recovery and recovered model structure as seen in Figures 5a-c.
@@ 5.2 PMd Dataset
We next apply our method to recordings from monkey dorsal premotor cortex (PMd) during sequential reaches (Figure 4a), in which trial-wise movement directions are continuously distributed over the circle. Full experimental details are in (Lawlor et al., 2018) and the dataset is provided in (Perich et al., 2018). The trials’ reach direction spans the full polar range rather than clearly separated directions as in the Area2 dataset. For the PMd dataset, we similarly extract movement-related activity by taking a fixed window (100-100 to +500+500 ms) around the movement onset for each trial. We analyze PMd activity with the $5$-dimensional kinematic as inputs, i.e., $(\text{x\mbox{-}velocity},\text{y\mbox{-}velocity},\text{x\mbox{-}acceleration},\text{y\mbox{-}acceleration},\text{speed})$, and the first $16$ PCs as outputs. Following the Area2 protocol, we train MoLDS across hyperparameters on the PMd training split and select the model by validation criteria (NLL/RMSE/BIC). Figure 6a reports validation results where the $4$-component MoLDS trained with the Tensor-EM method has better performance. Figure 6b shows test one-step-ahead reconstructions of the corresponding component from this MoLDS fit.
Figure 6: MoLDS application on PMd Dataset: (a) Validation metrics of MoLDS for different $K$ (NLL, RMSE, BIC). (b) One-step prediction on an example neural trial using the corresponding component from the validation-chosen MoLDS. (c) Angle bins: dominant component; PD arrows (responsibility-weighted). (d) Impulse response magnitude of Tensor-EM MoLDS components. (e) Tensor-EM MoLDS component usage fractions on held-out test trials.
To accommodate continuously distributed reach angles, we uniformly binned trials into $12$ angular bins based on their reach angles. For each bin, we assigned a dominant component based on the highest mean responsibility. Figure 6c shows these bins colored by their dominant LDS component and overlays responsibility-weighted preferred-direction (PD) arrows for each component, where we also compare Tensor-EM and Random-EM MoLDS results. Figure 6d plots impulse-response magnitude across lags of each component, revealing component-specific gains and temporal decay; Appendix E.2 (Figure 14) further shows the decomposed responses across input channels. Figure 6e reports Tensor-EM MoLDS component usage fractions on held-out test trials. Overall, the mixture components of the PMd dataset specialize in different movement directions and exhibit distinct dynamical response profiles.